### v2

# Hype-Adjusted Probability Measure for NLP Stock Return Forecasting

Abstract
This manuscript introduces the Hype-Adjusted Probability Measure developed in the context of a new Natural Language Processing (NLP) approach for stock return and volatility forecasting. A novel sentiment score equation is presented to capture component and memory effects and assign dynamic parameters, enhancing the impact of intraday news data on forecasting next-period volatility for selected U.S. semiconductor tickers. This approach integrates machine learning techniques to analyze and improve the predictive value of news. Building on the research of Geman et al [deveikyte2022], this work improves forecast accuracy by addressing news bias, memory, and weight, and incorporating shifts in sentiment direction. Finally, we propose the Hype-Adjusted Probability Measure, proving its existence and uniqueness, and discuss its theoretical applications in finance for NLP-based stock return forecasting, outlining future research pathways inspired by its concepts.
Keywords: Hype-adjusted Probability Measure, Natural Language Processing, Sentiment Analysis, Market Volatility Forecast, Semiconductor Industry
1 Introduction
Attention generates capital; capital shifts draw attention, and hype serves as the catalyst.
The inherent randomness of financial markets implies that precise predictions of future volatility and returns for any asset or its derivatives are almost surely impossible. Many pioneers have built sophisticated analytical and numerical instruments to evaluate future prices under strict assumptions, such as the well-known Black-Scholes model for an ideal risk-neutral world [BlackScholes1973]. Extensive research has shown that calibration and theoretical assumptions are essential to enhance model performance.
Sentiment has long been recognized as a key driver in financial markets, influencing both trading behavior and asset prices. Previous studies have examined the theoretical and empirical role of sentiment in market dynamics, highlighting its impact on investors’ decision-making and trading patterns [hirshleifer2001role, zhong2018sentiment]. Furthermore, research shows a significant relationship between investor sentiment and stock returns, providing a foundation for sentiment-based trading strategies [baker2006investor].
This manuscript contains key observations from “A Sentiment Analysis Approach to the Prediction of Market Volatility", where Latent Dirichlet Allocation (LDA) is applied to forecast market price and volatility [deveikyte2022]. The authors suggested an advanced method to compute the sentiment score and introduced how to forecast the magnitudes of the next day’s market return and volatility, in addition to predicting future stock return and volatility change.
We choose semiconductors as the primary sector for analyzing news and market data. We first review previous works, and propose the market intuition of the “3 weights" and numerical improvements based on the new sentiment score equation and NLP approach through the improved forecast results using ML.
We first review the past literature on NLP and Change of Measure in section 2. A detailed overview of data and methods is presented in section 3 which includes the developments from previous sentiment equation to more sophisticated ones along with enhanced NLP forecasting methods. Building upon these foundations, we introduce a novel probability measure, the Hype-Adjusted Probability Measure, $\mathbb{P}^{a}$, in section 4, designed to capture the effects of market "hype." Lastly, the results and future work are discussed in section 5.
2 Literature Review
2.1 Literature Review for NLP
Sentiment analysis, a subfield of Natural Language Processing (NLP), focuses on quantifying the emotional tone and intent conveyed in textual data. NLP itself is a rapidly growing area of machine learning (ML) that enables computers to process and understand human language through algorithms and statistical models [jurafsky2009speech]. Within NLP, sentiment analysis applies techniques to assess positive, negative, or neutral sentiment in data, often used for applications like market forecasting and consumer behavior analysis [liu2012sentiment].
In this context, sentiment is defined as the emotional polarity associated with a piece of text, often derived from news articles, tweets, or other media. It is typically measured using computational methods such as dictionary-based approaches or machine learning models. For example, Vader, a lexicon-based sentiment analysis tool, evaluates sentiment scores ranging from extremely negative to extremely positive [hutto2014vader].
In their previous research, Hélyette Geman and her colleagues identified a significant negative correlation between daily negative sentiment and observed market volatility and developed a sentiment-based model for predicting directional volatility and market returns [deveikyte2022]. Building on their findings, we aimed to enhance the predictive relationship between news sentiment and stock performance by refining the methods used to calculate key financial metrics.
The return on day $t$ is defined as the logarithmic change in the closing price from day $t-1$, expressed as:
$r_{t}=\log\frac{\text{CLOSE}_{t}}{\text{CLOSE}_{t-1}}$ (1)
Additionally, we adopted their method for defining the volatility, $\sigma$, which is calculated as:
$\sigma=\sqrt{\frac{1}{N}\sum_{t+1}^{N}(r_{t}-\bar{r})^{2}\cdot\sqrt{252}}$ (2)
Notice, for the programming and machine learning components of the manuscript, we selected a rolling window of 5 days to account for the amount of a regular trading week (without holiday breaks).
These metrics serve as the foundation for analyzing the connection between sentiment data and stock market behavior.
We adopted the Vader Sentiment engine to help translate multi-language texts into scores. It classifies sentiment scores on a scale as introduced by Clayton J. Hutto and Eric Gilbert [hutto2014vader]:
  1. 1.
    Extremely negative: -4
  2. 2.
    Neutral: 0
  3. 3.
    Extremely positive: +4
The base sentiment function introduced in [deveikyte2022] is defined as the fraction of the difference between the number of positive and negative sentiment tokens, scaled by the sum of positive, neutral, and adjusted negative cases:
$Sent_{d}=\frac{N_{d}(\text{pos})-N_{d}(\text{neg})}{N_{d}(\text{pos})+N_{d}(\text{neutral})-N_{d}(\text{neg})+3}$ (3)
One significant limitation of the prior sentiment equation is that it only counts the number of positive, negative, and neutral news. A news of sentiment score $-3$ should carry a bigger weight than a $-0.1$ score.
As a part of his master’s thesis, Zheng Cao, lead author of this paper, proposed a modified approach to sentiment scoring in April 2024 at Johns Hopkins University [CaoGemanThesis]. First, an interval of $(-0.05,+0.05)$ was chosen to broaden the range of sentiment scores classified as neutral, setting all scores within this domain to 00. Additionally, sentiment scores of extreme values, such as +4, are given higher weights than minimal positive values (e.g., +0.1), addressing the limitation in the original model where they had equal weights.
$Sent_{d}=\frac{1}{n}\sum_{i=1}^{n}SentScore_{i,d}$ (4)
Cao further incorporated a time-memory weight effect in sentiment calculations, represented as follows:
$Sent_{d}=$ $\omega_{\text{time weight},d}\frac{1}{n}\sum_{i=1}^{n}SentScore_{i,d}$ (5)
$+\omega_{\text{time weight},d-1}Sent_{d-1}$ (6)
These equations serve as the foundation for the developments of more sophisticated sentiment equations and the proposal of the hype-adjusted probability measure in the later sections.
2.2 Literature Review for Change of Measure
We draw inspiration for the change of measure from seminal works such as Harrison and Kreps’ 1979 introduction of the risk-adjusted measure for option pricing, which laid the foundation for modern financial mathematics by connecting the physical measure to pricing probabilities [HarrisonKrep1979]. Their work established the most fundamental cornerstone in this field, focusing primarily on stock returns rather than option pricing.
Building on this, Geman’s groundbreaking work in her 1989 dissertation, later published in “Changes of Numéraire, Changes of Probability Measure and Option Pricing" [Geman1995], extended these ideas to option pricing with stochastic interest rates by showing that changes in probability measures are intrinsically linked to changes in the Nnméraire.
More recently, Glasserman, Mamaysky, and Qin (2023) introduced the concept of entropy-based news novelty as a predictor of negative stock market returns and macroeconomic outcomes [glasserman2023new]. Their work leverages changes in the distribution of news text, quantified through entropy measures derived from recurrent neural networks. This innovative approach highlights the role of shifting language patterns in financial forecasting, offering a complementary perspective to traditional probability measure transformations.
In our work, while we acknowledge these advancements, we focus specifically on the representation of stock returns under measure changes. Although Geman et al.’s remarkable insights on measure changes for option pricing provide a rich theoretical basis, we intentionally leave the detailed discussion of these concepts for future research. The present paper aims to introduce the change of measure framework in the context of stock returns and sentiment-driven probability measures, laying the groundwork for further exploration in option pricing and beyond.
3 Data and Methods
In this study, ‘news’ refers to textual data from verified financial sources reporting on market trends, corporate developments, and stock performance.
Primary data sources for this study were obtained from LSEG and the Eikon API [LSEG], which provided access to over half a million news articles related to 30 semiconductor stocks. These data sources offered comprehensive coverage of market events, corporate developments, and sentiment indicators.
To mitigate potential biases in the collected data, particularly those arising from media overrepresentation or underrepresentation of specific events, we implemented two primary adjustments:
  1. 1.
    Assigning Weights to News Sources: Each news source was assigned a weight to achieve a more objective sentiment score adjustment. The adjusted sentiment score was calculated as:
    $\text{sentiment}_{\text{adj}}=\alpha\cdot\text{sentiment}+\beta$ (7)
    where $\alpha$ represents the weighting factor applied to a news source, and $\beta$ corrects for inherent biases.
  2. 2.
    Adjusting for Over- or Under-Reported Events: To address the uneven representation of events in the dataset, we applied two key measures:
    • Removed duplicate or near-identical articles to prevent overrepresentation of certain events.
    • Amplified the impact of under-reported events by assigning higher weights or using clustering techniques to identify and include diverse event representations.
3.1 Data and Naive Model Results
Appendix A presents a sample table of 30 tickers selected from the iShares Semiconductor ETF (SOXX). Note This ticker weight table is adjusted by removing 2 values: CME E-MINI S&P500-TECHNOLOGY SECTOR INDEX FUTURE SEP 2024,$0.15247\%$ and CME INDEX and OPTIONS MARKET E-MINI RUSSELL 2000 INDEX FUTURE SEP 2024, $0.15247\%$. The reason is that these two entries do not represent any actual company; they are financial derivatives or futures contracts tied to broader indices rather than individual equities, and therefore their inclusion would distort the sector-specific analysis focused on individual component stocks.
We first assemble the results using the previous sentiment methods on the semiconductor data:
  1. 1.
    the “daily weighted average" sentiment score: computed by the aggregated news data based on individual tickers, and adjust the processed sentiment score by individual tickers’ weight from the overall SOXX sector (excluding the 2 Future ETFs removed)
  2. 2.
    the “overall daily average" sentiment score: searched by ticker name, without being adjusted by individual weights
  3. 3.
    the “overall daily average semi title" sentiment score: searched by the topic of SOXX sector, without being adjusted by individual weights
Each data set generates a confusion matrix report through a simple Linear Discriminant Analysis algorithm [balakrishnama1998lda]. Linear Discriminant Analysis (LDA) is a commonly used dimensionality reduction technique that is effective for classification. It projects data onto a lower-dimensional space while maximizing separation between classes by modeling inter-class differences using the mean and variance within each class. Unlike Principal Component Analysis (PCA), which maximizes variance, LDA prioritizes class separability, making it especially useful in supervised learning with labeled data.
For the “daily weighted average" sentiment score:
Class Precision Recall F1-Score Support
0 0.65 0.73 0.69 33
1 0.62 0.54 0.58 28
Accuracy 0.64
Macro Avg 0.64 0.63 0.63 61
Weighted Avg 0.64 0.64 0.64 61
Table 1: Classification Report for Best Volatility Direction Model
For the second, the “overall daily average" sentiment score:
Class Precision Recall F1-Score Support
0 0.58 0.91 0.71 32
1 0.73 0.28 0.40 29
Accuracy 0.61
Macro Avg 0.65 0.59 0.55 61
Weighted Avg 0.65 0.61 0.56 61
Table 2: Classification Report for Best Volatility Direction Model
For the third data input, the “overall daily average title" sentiment score, the result is
Class Precision Recall F1-Score Support
0 0.66 0.87 0.75 31
1 0.80 0.53 0.64 30
Accuracy 0.70
Macro Avg 0.73 0.70 0.70 61
Weighted Avg 0.73 0.70 0.70 61
Table 3: Classification Report for Best Volatility Direction Model
The latter sections present a more sophisticated equation of processing the sentiment score, with supporting results provided.
Figure 1 presents a sample distribution of calculated sentiment scores, showing a notable negative skew.
Figure 1: Distribution of Sentiment Scores
This serves as a motivation for adjusting bias weights, as discussed in Section 3.2, which account for market ‘hypes’ and form the basis for the proposed hype-adjusted probability measure. Bias weights are adjusted to correct for distortions such as excessive coverage of certain stocks and underrepresentation of others. These adjustments ensure a more objective and balanced assessment of market dynamics. Without such corrections, the analysis may disproportionately emphasize certain stocks or news sources, resulting in skewed forecasts or inaccurate conclusions.
3.2 Bias, Memory, and Weights
This section explores the foundational elements - bias, memory, and weights - that underpin the construction of a refined sentiment equation, serving as the basis for the proposed hype-adjusted probability measure. The goal is to systematically address and correct biases arising from media attention imbalances and market weighting discrepancies while incorporating the dynamic influence of past sentiment.
3.2.1 News Bias
News bias refers to the disparity in media coverage that individual stocks or sectors receive, which may not align with their actual market significance. For instance, major companies like Nvidia tend to dominate media attention, potentially distorting the sentiment analysis for the entire sector.
The ticker news count weight represents the proportion of total news articles attributed to a specific ticker compared to all tickers within the sector. This metric quantifies the relative attention a ticker receives from the media and audience in comparison to other tickers in the same domain or sector. Formally, for ticker $i$:
$\text{Ticker News Count Weight}_{i}=\frac{\text{News Count}_{i}}{\sum_{j=1}^{n}\text{News Count}_{j}},$ (8)
where $n$ is the total number of tickers in the sector. This weight reflects how much media coverage a particular ticker receives relative to others.
The new bias arises from the gap between the Ticker News Count Weight and the actual market weight of a ticker. Mathematically:
$\text{Bias}_{i}=\text{Ticker News Count Weight}_{i}-\text{Market Weight}_{i},$ (9)
where:
  • A positive bias ($\text{Bias}_{i}>0$) indicates over-representation in the news.
  • A negative bias ($\text{Bias}_{i}<0$) suggests under-representation in the news.
Adjusting for this bias is essential to ensure that the sentiment analysis accurately reflects market dynamics rather than being skewed by excessive coverage or neglect of specific tickers. Without such adjustments:
  • Over-represented tickers (e.g., Nvidia) may disproportionately influence sector sentiment, leading to inflated predictions.
  • Under-represented tickers may not sufficiently contribute to the overall sentiment, resulting in overlooked trends.
By integrating both Ticker News Count Weight and Market Weight into the sentiment equation, we aim to balance these disparities and improve the accuracy of market sentiment models.
We examine news bias within the iShares Semiconductor ETF (SOXX) of the U.S. stock market by analyzing how individual asset components contribute to the overall index. Several major companies, such as Nvidia, dominate the sector, with Nvidia representing over $8\%$ of the market. However, over the past 15 months (up to mid-July 2024), Nvidia has received disproportionately high news coverage of $24.52\%$ compared to other companies in the sector. Considering 30 companies from the SOXX, we assign weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%. Please refer to Appendix A for more details.
Figure 2: Actual Linear Relationship of Ticker News Count Weight vs Capital Worth Weight
A zoom on the left bottom side of the plot is provided by filtering out NVDA.QQ and INTC.QQ from the ticker list.
Figure 3: Actual Linear Relationship of Ticker News Count Weight vs Capital Worth Weight, Excluding NVDA and INTC
We wish to provide some sophisticated method to change the relationship from the first plot Figure 2 to the second in Figure 4.
Figure 4: Ideal Linear Relationship of Ticker News Count Rank vs Capital Worth Rank
Further data and analysis are provided in the Intuition of the New Measure subsection.
To address news bias, we consider two key factors:
  1. 1.
    Event Exaggeration: Some media outlets tend to sensationalize events, which can disproportionately impact investor decisions and market reactions.
  2. 2.
    Selective Focus on High-Profile Companies: Media often emphasizes a few prominent companies, amplifying their perceived influence and neglecting others.
We propose two approaches to address these biases:
  1. 1.
    Assigning Weights to News Sources: Each news source is assigned a weight to adjust sentiment scores more objectively. For example:
    $\text{sentiment}_{adj}=\alpha\cdot\text{sentiment}+\beta$ (10)
    where $\alpha$ controls the sentiment weight from the source, and $\beta$ corrects for inherent bias.
  2. 2.
    Adjusting for over- or under-reporting: To correct for uneven media coverage, we adjust the volume of reported news by removing duplicates or overly similar articles.
3.2.2 Memory and Weight
Quantifying the bias of news sources and optimizing compound sentiment scores are key to improving sentiment-based forecasting models. We propose a weighted approach that integrates current sentiment with past sentiment, adjusting for biases:
$Sent_{d}=$ $\ \omega_{\text{today weight}}\sum_{i=1}^{n}\omega_{\text{news bias}}\omega_{\text{component weight}}\,SentScore_{i,d}$ (11)
$+\omega_{\text{past weight}}Sent_{d-1}$ (12)
Here:
  • $d$: Represents today or the current period of time. $d-1$ refers to yesterday or the previous period of time.
  • $n$: Denotes the total number of tickers studied.
This equation includes:
  • Present sentiment: A weighted sum of sentiment scores adjusted for news bias ($\omega_{\text{news bias}}$) and component weights ($\omega_{\text{component weight}}$).
  • Past sentiment: The sentiment from the previous day ($Sent_{d-1}$) is included, weighted by its relevance over time.
By incorporating these weights, the model ensures that the sentiment score reflects both the immediate impact of current news and the lingering effects of past sentiment, thereby improving the accuracy of market forecasts.
The primary goal is to determine the extent to which news sources exaggerate or understate events and how this impacts market sentiment. By incorporating bias weights, memory effects, and component-specific adjustments, the model aims to better reflect true market sentiment, minimizing the effects of biased reporting. Future improvements could involve more advanced bias detection algorithms and real-time integration of data from platforms like Eikon.
Future work can be to continue improving the quantifying method for the news bias.
3.2.3 Recursive Method
We propose new parameters for calculating the daily compound sentiment score. Rather than using daily averages as a benchmark for $Sent_{d}$ (sentiment score on the date or time $d$), we assign weights to each component (ticker) based on its sector weight:
$Sent_{d}=$ $\omega_{\text{today weight}}\sum_{i=1}^{n}\omega_{\text{componentweight}}SentScore_{i,d}$ (13)
$+\omega_{\text{past weight}}Sent_{d-1}$ (14)
Using the semiconductor sector as an example, we analyze how individual asset components contribute to the overall index. Considering 30 companies from the SOXX, we assign weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%.
$\sum_{n=0}^{29}\omega_{n}=1$ (15)
We introduce a novel approach to account for time-memory effects in market reactions. Historical news sentiment impacts future market performance, with older news exerting diminishing influence. For instance, the collective sentiment on a Monday may be shaped not only by weekend news but also by earlier events. We propose weighting algorithms to capture these decay and lagging effects, ensuring older events carry less influence while accounting for delayed impacts across different assets.
We propose that shifts in sentiment direction affect the relevance of historical data. This is incorporated into the construction of indicator functions in the new sentiment score equation, adjusting the weight of past data based on directional changes.
3.2.4 Summary: The 3 Weights
We examine the impact of three types of weights on market sentiment analysis:
1. Market Component Weights vs News Coverage:
Certain stocks receive disproportionate media coverage compared to their actual weight in the market or sector. For instance, Nvidia may garner excessive news coverage relative to its actual weighting in the semiconductor sector.
To address this imbalance, we propose a solution that adjusts the contribution of individual tickers’ sentiment to the overall sentiment of the industry or sector. This approach aims to mitigate the disproportionate influence of certain stocks, progressing from the current scenario (top figure) to a more balanced perspective (bottom figure).
Currently, the approach is to train the algorithm to learn the best weight for each individual component
2. News Source Bias Weights:
We recognize that news reports are rarely perfectly objective. Different news sources have biases that can affect sentiment. For example, CNN might exhibit a liberal bias and present a relatively positive outlook on green energy, whereas FOX may be more critical of gun control policies. Similarly, the BBC may adopt a more critical stance on the Chinese market compared to state-owned Chinese media, with biases shaped by the source’s audience and language.
To correct for these biases, we propose assigning weights to different news sources based on the degree of bias and applying these weights when analyzing sentiment data.
To quantify the news bias, we plan to use data from different media sources and predict the results separately. We assume that a more accurate forecast comes from a more objective data source. And we will apply the learned degree of bias to assign new weights to compute the adjusted sentiments.
3. Weights of Past News Data (Memory):
We have already demonstrated that adjusting sentiment using indicator functions, coupled with machine learning techniques to optimize parameters, can enhance the accuracy of market forecasts.
Building on this, we aim to derive a new score equation that captures how the weighting of past news data (memory) influences current prices. This weighting function will help us better understand how historical sentiment data affects current market dynamics.
In summary, we focus on evaluating and optimizing the “3 weights" to enhance the sentiment score equation and NLP approach. These enhancements ensure a balanced and accurate representation of market dynamics by correcting for media attention imbalances, accounting for source biases, and incorporating memory effects. This foundational framework not only refines sentiment modeling but also provides the necessary groundwork for the development of an improved sentiment score equation. In the next section, we extend these concepts to introduce a more robust equation that integrates component weighting and historical memory effects for greater predictive accuracy.
3.3 Improved Sentiment Score Equation
This section presents an enhanced sentiment score equation that incorporates the weighting of individual components within a portfolio and accounts for historical memory effects from past news data.
Notations
  1. $1$
    $Sent_{d}$: The compound daily sentiment score of an underlying asset portfolio, calculated without considering historical data.
  2. $2$
    $SentScore_{i,d}$: The average sentiment score for news related to ticker $i$ on date $d$, based on a total of $k$ news articles.
  3. $3$
    $SentAll_{d}$: The compound daily sentiment score incorporating the cumulative influence of historical data.
$Sent_{d}=$ $\ \sum_{i=1}^{n}\omega_{i}\,SentScore_{i,d}$ (16)
where $n$ represents the total number of tickers in the portfolio.
$SentScore_{i,d}=\frac{1}{k}\ \sum_{j=1}^{k}S_{i,d,j}$ (17)
where $SentScore_{i,d}$ denotes the average sentiment score for ticker $i$ on date $d$, calculated from a total of $k$ news articles.
We incorporate only the most recent significant sentiment shift event. If multiple shifts have occurred but the threshold for disregarding prior data has not been met, only the most recent shift is considered, and sentiments from earlier dates are neglected.
$SentAll_{d}=$ $\sum_{i=1}^{k}\mathbbm{1}_{Sent_{d},Sent_{d-1}}\quad\omega_{\text{weight}_{i,d}}SentScore_{i,d}$ (18)
$=$ $\omega_{\text{today weight}}Sent_{d}$
$-\mathbbm{1}_{Sent_{d}<0\ \&\ Sent_{d-1}>0}\quad\omega_{\text{past weight 1}}\,SentAll_{d-1}$
$+\mathbbm{1}_{Sent_{d}>0\ \&\ Sent_{d-1}<0}\quad\omega_{\text{past weight 2}}\,SentAll_{d-1}$
$+\mathbbm{1}_{Sent_{d}\times Sent_{d-1}\geq 0}\quad\omega_{\text{past weight 3}}\,SentAll_{d-1}$
We consider the influence of historical sentiment. The $SentAll_{d}$ variable represents the compound daily sentiment score, taking into account the cumulative influence of historical data. Each term in this equation has the following interpretation:
We categorize sentiment changes into three cases to assess the role of historical data in the overall sentiment function, based on indicator functions:
  1. 1.
    $\mathbbm{1}_{Sent_{d}<0\ \&\ Sent_{d-1}>0}$ : when the daily sentiment shifts from positive to negative.
  2. 2.
    $\mathbbm{1}_{Sent_{d}>0\ \&\ Sent_{d-1}<0}$: when the daily sentiment shifts from negative to positive.
  3. 3.
    $\mathbbm{1}_{Sent_{d}\times Sent_{d-1}\geq 0}$ : when the daily sentiment directions stay unchanged
When the indicator $\mathbbm{1}_{Sent_{d}<0\ \&\ Sent_{d-1}>0}$ is met, higher volatility is likely to occur as compared to the condition $\mathbbm{1}_{Sent_{d}>0\ \&\ Sent_{d-1}<0}.$
The real-world implication leads us to hypothesize that the more positive the overall sentiment is (the greater $SentAll_{d-1}$ is), the more impacts it will cause from negative news occurrences to traders’ panic sales and the drop in market prices (the smaller $SentAll_{d}$ becomes, and therefore introducing the negative sign before $\mathbbm{1}_{Sent_{d}<0\ \&\ Sent_{d-1}>0}$.
We hypothesize that when sentiment shifts from positive to negative, the impact becomes more significant as the magnitude of the change increases. This would directly correlate with the weight we aim to introduce here and may warrant further study.
where
  • $\omega_{\text{time memory},d}$ depends on how many days ahead, and whether it is a trading day or not,
  • $\omega_{\text{component weight}}$ depends on the individual market weight of the ticket in its sector.
Weights below $0.005$ are disregarded. For instance, after multiple recursive applications, if $\omega_{\text{time weight},d-10}=0.003<0.005$, data for dates on or before $d-10$ are excluded from consideration.
3.3.1 Continuous Sentiment Equation
To convert the discrete sentiment score equation into a continuous form using integrals, we replace summations with integrals and assume that $Sent_{d}$, $Sent_{d-1}$, and $SentAll_{d}$ represent continuous sentiment functions over time, $t$.
The discrete sum:
$SentAll_{d}=\sum_{i=1}^{k}\mathbbm{1}_{Sent_{d},Sent_{d-1}}\omega_{\text{weight}_{i,d}}SentScore_{i,d}$
can be replaced by an integral form over a continuous time period:
[…]
We now transform each case in the discrete equation to its continuous form:
  1. 1.
    Case 1: When $Sent(t)<0$ and $Sent(t-\Delta t)>0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
  2. 2.
    Case 2: When $Sent(t)>0$ and $Sent(t-\Delta t)<0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
  3. 3.
    Case 3: When $Sent(t)\times Sent(t-\Delta t)\geq 0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
These integrals represent the continuous version of how historical sentiment influences the current sentiment based on the indicator functions $I$, which trigger depending on the changes in sentiment over time. Each term involves weighting past sentiment contributions and the present day’s sentiment to compute the total sentiment.
Note that the period does not need to be daily; the gap interval can be shorter for intraday analysis, such as every hour, allowing for finer granularity in processing and analysis.
In the following section, we apply the enhanced sentiment score equation within an updated NLP framework, testing its predictive power using various algorithms and assessing its impact on market forecast accuracy.
3.4 New NLP Approach and Forecast Results
This section summarizes the machine learning (ML) methodologies, procedures, and forecast results used in the new NLP approach, which is based on the enhanced sentiment score equation.
3.4.1 Methodologies
The primary method introduced in the earlier paper relies on LDA. In this study, however, we use unsupervised ML models as benchmarks to compare the performance of enhanced computational and optimization techniques.
Our objectives are twofold: first, to improve the forecasting accuracy of the original model, and second, to demonstrate the benefits of incorporating arithmetic computations and optimizations beyond standard LDA.
Figure 5: Algorithm Flow Chart
The optimization algorithms used include models such as LDA, Logistic Regression, Ordinary Least Squares (OLS) regression. The primary focus of this paper is not to fine-tune machine learning models but to demonstrate the improvements gained from the new sentiment score data and procedure.
3.4.2 Machine Learning Procedures
We employed several machine learning approaches to predict stock market movements based on sentiment data. The sentiment data was first processed by converting the date index into a standardized date format. This data was then merged with stock market data, including daily market returns and volatility metrics.
The methodology involved iterating through various sentiment indicators and performing data splits into training, validation, and test sets. For each indicator, we ran the model up to 1,000 times with different random states to ensure robust results.
Ordinary Least Squares (OLS) regression was used to quantify the relationship between sentiment scores (independent variable $X$) and daily market return or volatility (dependent variable $Y$). The regression provides an R-squared value to measure the proportion of variance in $Y$ explained by $X$. This allows us to evaluate the predictive power of sentiment scores in explaining market dynamics.
Logistic regression was employed to predict the direction of daily market returns (dependent variable $Y$) based on the compound sentiment score (independent variable $X$). Here, $Y$ was a binary variable indicating positive or negative market movements. The accuracy of the logistic regression model served as a key metric for assessing its performance.
The optimal models were selected based on the highest validation accuracy (for logistic regression) and R-squared values (for OLS regression) across iterations. These models were subsequently evaluated on the test dataset, with metrics including accuracy, precision, recall, and confusion matrices. The results highlighted the most effective sentiment indicators for predicting market movements, offering insights into the relationship between sentiment data and market behavior.
Sample table showing modified sentiment scores based on different parameters (for equation 18) from the new equation:
Table 4: Sample Sentiment Scores with Selected Parameters
Date Base Sentiment Parameters A Parameters B Parameters C
2023-04-27 0.043192 0.043192 0.043192 0.043192
2023-04-28 0.186383 0.186383 0.207979 0.229574
2023-04-29 0.040962 0.040962 0.144951 0.270536
2023-04-30 -0.036497 -0.036497 -0.036497 -0.036497
2023-05-01 0.078847 0.078847 0.078847 0.078847
Each column represents the adjusted sentiment scores based on different sets of parameters.
The model results are analyzed in Section 5.1.
4 Hype-Adjusted Probability Measure
4.1 Intuition of the New Measure
While we previously proposed a sentiment equation in discrete form, we now extend it to a continuous framework to capture market dynamics with greater precision. The limitations of discrete methods prompt us to provide mathematical justifications alongside market implications, addressing parameters like the “3 weights" introduced earlier.
In the context of financial markets, hype refers to the amplification of attention or sentiment around a particular stock, sector, or event that exceeds its fundamental or intrinsic importance. Hype is often fueled by disproportionate media coverage, speculative behavior, or investor overreaction to news.
A hype scenario is identified when there is a measurable and disproportionate increase in examples like:
  1. 1.
    Media Coverage: A significant spike in the volume of news articles, social media mentions, or other sources of information about a specific stock or sector compared to its baseline or relative importance.
  2. 2.
    Market Reactions: Corresponding anomalies in price movement and volatility, such as sharp increases or dramatic swings, that deviate from historical patterns.
  3. 3.
    Imbalance in Representation: Evidence of over or under representation in news coverage compared to the stock’s weight in its sector (e.g., market capitalization).
These indicators form the foundation for quantifying and incorporating hype into the proposed probability measure. A real-world case of Nvidia’s hype is examined later in this section.
Recall the “3 weights":
  • Market Component Weights vs news Coverage"
  • News Source Bias Weights
  • Weights of Past News Data (Memory)
These parameters capture the influence of market components, news biases, and historical data retention (memory) on sentiment and subsequently on the market forecast.
Also, recall that we have shown from figures that a slight negative skew of market news sentiments is observed and there exists an imbalance of news report counts vs. component weights of Nvidia among the semiconductor sector tickers.
In addition, the news source bias weights are currently being investigated for further research. For example, in America, major news media have political leanings, CNN is more left-wing and promotes clean energy and gun control, while Republican-favored Fox argues for the opposite tone. We wish to quantify the bias instead of relying solely on polls and surveys.
To illustrate the concept of market ‘hype’, we examine Nvidia, the most trending AI company as a case study for summer 2024.
Figure 6: Number of News for Nvidia Q1 Earning Report on May 22, 2024, versus tickers in SOXX
In figure 6, the blue line indicates the news counts on Nvidia from May 10 to June 9, 2024, while the red dashed line indicates May 22, 2024, when Nvidia’s Q1 earnings report was released. The orange trajectory describes the total number of news collected for all tickers in the SOXX sector. We notice a rapid increase in market hype centering around the report release time.
Figure 7: Price and Volatility Trajectories for Nvidia Q1 Earning Report on May 22, 2024
Figure 7 presents the historical movements of Price and Volatility based on the market close data for the same time frame of Figure 6, it is on observation of a dramatic increase in both the close price and the close price volatility.
Notice, while we do not assert a direct positive correlation between news count increases and price or volatility changes, these observations suggest that market hype can significantly impact these variables, which inspires the proposal of the hype-adjusted probability measure, $\mathbb{P}^{a}$.
4.2 Hype-Adjusted Probability Theory
Based on the intuitions, we define the hype-adjusted probability measure, $\mathbb{P}^{a}$, to account for market sentiment and the role of news weight biases.
We choose volatility as the target of prediction, which aligns with the observation results from sections 3.4 and 4.1. We also discuss the relationship between the hype-adjusted probability measure and the volatility and how to apply this for market price forecasting.
Definition: Hype-adjusted Probability Measure
The hype-adjusted probability measure, denoted $\mathbb{P}^{a}$, is a probability measure that incorporates sentiment biases by adjusting the physical measure $\mathbb{P}$. The probability space for hype-adjusted probability measure is a triplet $(\Omega,\mathcal{F},\mathbb{P}^{a})$. Under $\mathbb{P}^{a}$, the effects of news weight, bias from media sources, and sentiment shifts are inherently accounted for, allowing a more accurate representation of expected market behaviors influenced by hype.
4.2.1 Change from Physical to Hype-Adjusted Probability Measure
We get the inspiration of change of measure from the physical measure to Harrison and Kreps’ risk-adjusted measure (1979) and to Geman’s Forward measure (1989) as introduced in the literature review section 2.2.
For simplicity, we illustrate the change of measure by considering only the "Capital vs. News Counts" factor of the “3 weights."
  • Let $L_{Cap}(\omega)$ denote the capital weight of a given event (ticker).
  • Let $L_{News}(\omega)$ denote the news weight of a given ticker.
As shown in Appendix A, it is clear that:
$\sum L_{Cap}(\omega)=\sum L_{News}(\omega)=1$
Hence, we assign
$\mathbb{P}(\omega)=L_{News}(\omega)\quad\text{and}\quad\mathbb{P}^{a}(\omega)=L_{Cap}(\omega),$
which leads to:
$\sum\mathbb{P}(\omega)=\sum\mathbb{P}^{a}(\omega)=1.$
It follows almost surely that no weight is equal to zero. In other words:
$\text{Prob}(L_{Cap}(\omega)=0)=0\quad\text{and}\quad\text{Prob}(L_{News}(\omega)=0)=0.$
Therefore, by the Equivalent Measure Theorem, the two probability measures agree on which sets in $\mathbb{F}$ have probability zero. Intuitively, an neutral sentiment score ($=0$) under $\mathbb{P}$ is still neutral under $\mathbb{P}^{a}$.
The change of measure theorem states that, since $\mathbb{P}$ and $\mathbb{P}^{a}$ are equivalent, the Radon-Nikodym derivative
$\mathbb{Z}(\omega)=\frac{d\mathbb{P}(\omega)}{d\mathbb{P}^{a}(\omega)}$ (20)
exists. For any $X\in\mathcal{F}$, we have:
[…]
This adjustment allows us to account for market hype by converting the original news weight of a particular ticker to its correct proportion within the sector.
Remark - Sentiment Adjustment:
For the hype-adjusted probability measure $\mathbb{P}^{a}$, the Radon-Nikodym derivative $\mathbb{Z}(\omega)$ adjusts the impact of news on tickers based on their level of media coverage. Specifically:
  • For over-hyped tickers (e.g., NVDA.QQ), the news count impact is discounted, represented by $\mathbb{Z}(\omega)<1$.
  • For under-represented tickers (e.g., MPWR.OQ), the news count impact is amplified, represented by $\mathbb{Z}(\omega)>1$.
This adjustment balances sentiment influence across both highly publicized and lesser-known tickers.
Recall, under $\mathbb{P}$, we have equation 19 as :
$\text{Sentiment}(t)$ $=SentAll(t)$ (21)
[…]
Under $\mathbb{P}^{a}$, we can rewrite equation 19 as:
[…]
where $\hat{\text{Sentiment}}(t)$ represents the adjusted sentiment score of a given period under $\mathbb{P}^{a}$.
We then obtain the expected sentiment score $\hat{\text{Sentiment}}(t)$ over time $t$ as:
[…]
where:
  • $\frac{d\mathbb{P}}{d\mathbb{P}^{a}}$: the Radon-Nikodym derivative, representing the adjustment from the original measure $\mathbb{P}$ to the hype-adjusted measure $\mathbb{P}^{a}$.
  • $\mathbbm{1}(Sent(s),Sent(s-\Delta s))$: an indicator function that activates based on sentiment shifts over the interval $[s,s-\Delta s]$.
  • $\omega_{\text{weight}}(s)$: the weighting function for sentiment impact at time $s$.
  • $SentScore(s)$: the sentiment score at time $s$.
  • $\mathcal{F}_{t}$: the filtration up to time $t$, representing the information available in the market.
Thus, the sentiment score $\hat{\text{Sentiment}}(t)$ under the hype-adjusted measure $\mathbb{P}^{a}$ is defined as:
Theorem 4.1 (Conditional Expectation of Sentiment).
The Conditional Expectation of Sentiment at time T under the hype-adjusted probability measure is:
[…]
where $\hat{SentScore}(T)$ represents the adjusted sentiment score under $\mathbb{P}^{a}$.
In this form, the adjusted measure $\mathbb{P}^{a}$ captures market sentiment dynamics directly through the sentiment score $\text{SentScore}(s)$ alone. The absence of explicit weighting and indicator functions in Equation 26 highlights that $\mathbb{P}^{a}$ automatically adjusts for sentiment influence and bias over time, streamlining the overall computation.
We propose:
Theorem 4.2 ($\mathbb{P}^{a}$ martingale).
There exits a $\mathbb{P}^{a}$ martingale, such that:
$\mathbb{P}^{a}\sim\mathbb{P}$
and
$\mathbb{E}^{\mathbb{P}^{a}}\left[\frac{\partial\hat{\text{Sentiment}}(s)}{\partial\text{Volatility(s)}}\mid\mathcal{F}_{t}\right]=\frac{\partial\hat{\text{Sentiment}}(t)}{\partial\text{Volatility(t)}},\quad\forall s>t,$ (27)
where $\sigma(s)$ is the volatility of underlying security prices.
4.2.2 Hype-Adjusted Measure Under Discrete Binomial Tree Model
The discrete binomial tree model provides a clear framework to illustrate the concept of changing measures in a discrete setting. Traditionally, the binomial tree represents potential future movements of asset prices or volatility over a series of discrete time intervals. Here, we adapt this model to transition from the physical measure $\mathbb{P}$ to the hype-adjusted measure $\mathbb{P}^{a}$ by adjusting the probabilities associated with upward and downward volatility movements, incorporating the "Capital vs. News Counts" factor as weights.
Unlike a standard binomial asset pricing tree model, we have:
  • At each time step, the volatility can move up by a factor $u$ with probability $p$ or down by a factor $d$ with probability $1-p$.
  • The volatility at time $t+\Delta t$ is given by:
    $\sigma_{t+\Delta t}=\begin{cases}\sigma_{t}u&\text{with probability }p,\\ \sigma_{t}d&\text{with probability }1-p.\end{cases}$
To introduce the hype-adjusted measure, define:
  • $L_{Cap}$: the capital weight associated with each event (ticker).
  • $L_{News}$: the news weight assigned to each event (ticker).
Under the physical measure $\mathbb{P}$, the probabilities for an upward and downward movement are given by $p=L_{News}(u)$ and $1-p=L_{News}(d)$, respectively. The hype-adjusted measure $\mathbb{P}^{a}$ adjusts these probabilities as follows:
$p=L_{Cap}(u)\quad\text{and}\quad 1-p=L_{Cap}(d).$
This adjustment captures the impact of sentiment-driven volatility changes, with $L_{Cap}$ and $L_{News}$ modulating the probabilities under the hype-adjusted measure.
The Radon-Nikodym derivative $\mathbb{Z}(\omega)$ allows us to transition from $\mathbb{P}$ to $\mathbb{P}^{a}$. In the binomial tree context, this derivative can be expressed at each node as:
$\mathbb{Z}(u)=\frac{L_{Cap}(u)}{L_{News}(u)},\quad\mathbb{Z}(d)=\frac{L_{Cap}(d)}{L_{News}(d)}.$
These adjusted probabilities reflect how market sentiment biases (e.g., disproportionate media focus on certain tickers) are incorporated into the measure.
For volatility $\sigma_{t}$, the expected value at the next time step under $\mathbb{P}^{a}$ is:
$\mathbb{E}^{\mathbb{P}^{a}}[\sigma_{t+\Delta t}\mid\sigma_{t}]=\sigma_{t}\left(L_{Cap}(u)u+L_{Cap}(d)d\right).$
In contrast, under $\mathbb{P}$, it would be:
$\mathbb{E}^{\mathbb{P}}[\sigma_{t+\Delta t}\mid\sigma_{t}]=\sigma_{t}\left(L_{News}(u)u+L_{News}(d)d\right).$
This adjustment aligns the asset’s future volatility expectations with the corrected market sentiment weightings, effectively capturing hype adjustments in a discrete framework.
4.2.3 Sentiment-Driven Volatility and Hype-Adjusted Pricing
Building upon the discrete analysis, our work leverages established relationships between volatility, stock prices, and sentiment, as demonstrated in models such as the Black-Scholes-Merton (BSM) model. In the BSM framework, volatility is a critical parameter that directly influences option pricing. Traditionally, volatility is viewed as a measure of market risk and is calculated based on historical price movements. However, we have shown that sentiment shifts, particularly those driven by hype or disproportionate news coverage, can significantly impact volatility.
Theorem 4.3 (Sentiment-Driven Volatility Fluctuations).
Let $\sigma(t)$ denote the volatility of an underlying asset and $\text{Sentiment}(t)$ represent the sentiment score at time $t$. We assume that fluctuations in volatility are sentiment-driven. Empirical analysis suggests a correlation between changes in sentiment $d\text{Sentiment}(t)$ and adjustments in volatility $d\sigma(t)$, such that:
$d\sigma(t)\approx\alpha\,d\text{Sentiment}(t),$ (28)
where $\alpha$ is a sensitivity parameter derived from historical relationships between sentiment and volatility, as proposed in equation 27.
This extension allows for sentiment-adjusted volatility $\hat{\sigma}(t)$, defined as:
$\hat{\sigma}(t)=\sigma(t)+\alpha\,d\text{Sentiment}(t),$ (29)
which reflects sentiment-driven adjustments.
The sentiment-adjusted volatility $\hat{\sigma}(t)$ serves as a dynamic input that adapts to sentiment variations, providing a refined approach to volatility forecasting in sentiment-sensitive markets.
Theorem 4.4 (Sentiment-Adjusted Stochastic Price Dynamics).
Under the hype-adjusted probability measure $\mathbb{P}^{a}$, the distribution of price trajectories incorporates sentiment-driven adjustments. The stochastic differential equation governing the asset price $S_{t}$ is modified to reflect sentiment-adjusted volatility, as follows:
$\frac{dS_{t}}{S_{t}}=r\,dt+\hat{\sigma}(t)\,dW_{t}^{\mathbb{P}^{a}},$ (30)
where $dW_{t}^{\mathbb{P}^{a}}$ represents the Wiener process under the hype-adjusted measure $\mathbb{P}^{a}$, and $\hat{\sigma}(t)$ is the volatility adjusted for sentiment effects.
This modification results in a set of stochastic trajectories that align price forecasts with sentiment-influenced market expectations, enabling a more accurate representation of market dynamics in sentiment-sensitive contexts.
By integrating the hype-adjusted measure $\mathbb{P}^{a}$ and sentiment-sensitive adjustments to volatility, we enhance traditional pricing models to better reflect the influence of market hype. This framework not only updates the pricing formula but also provides a probabilistic model for sentiment-adjusted price paths. Such insights enable more accurate derivative pricing in sentiment-driven markets and support refined investment strategies during periods of heightened market hype.
For example, with $\sigma(t)$ now incorporating sentiment adjustments in equation 29, we can update the standard Black-Scholes pricing formula for a European call option $C(S,t)$. The sentiment-adjusted volatility $\hat{\sigma}(t)$ replaces the traditional volatility $\sigma$ in the BSM formula, producing a sentiment-enhanced price:
$C(S,t)=SN(d_{1})-Ke^{-r(T-t)}N(d_{2}),$
where:
$d_{1}$ $=\frac{\ln\left(\frac{S}{K}\right)+\left(r+\frac{{\hat{\sigma}(t)}^{2}}{2}\right)(T-t)}{\hat{\sigma}(t)\sqrt{T-t}},$
$d_{2}$ $=d_{1}-\hat{\sigma}(t)\sqrt{T-t},$
At time $t$, we have information granted by the filtration $F_{t}$. We may decide which volatility $\sigma$ model or method to obtain the stock price $S(T)$ at future time $T$ and compare the results.
[…]
For example, at time $t$, we can derive volatility using one of the following methods:
  • The implied volatility $\sigma_{\text{Imp}}$ from the Black-Scholes equation.
  • The adjusted volatility $\hat{\sigma}$, incorporating news-based sentiment adjustments.
These choices lead to various trading strategies, such as a variance swap.
In summary, this section defines $\mathbb{P}^{a}$, from the change of measure from $\mathbb{P}$ to the analysis under the discrete binomial model, and the discussion for sentiment-driven volatility and hype-adjusted pricing.
5 Results and Future Work
5.1 Model Results
We observed that the expected average difference between the original volatility and the adjusted one, based on the weights provided in the dataset, is approximately $-0.0068$ (or $-0.68\%$). Significant improvements were observed with the new NLP approach utilizing the enhanced sentiment scoring equation.
Figure 8: Model Validation Accuracy Comparison for Market Return and Volatility Forecasting
Figure 8 demonstrates the progression of accuracy across different models. For predicting market return direction, accuracy improved from $51.7\%$ in the baseline model to $70.0\%$ with a Na"ive model using adjusted scores, and further to $78.3\%$ with the best-optimized scores. Similarly, for volatility direction, accuracy increased from $53.8\%$ in the baseline model to $72.1\%$ and $75.0\%$ in the corresponding models. These substantial improvements validate the effectiveness of our framework.
The observed accuracy increase of $+8.3\%$ for market return direction and $+2.9\%$ for volatility direction highlights the refined precision achieved through optimized sentiment modeling. This improvement translates to more reliable predictions of market trends, a critical factor for decision-making in high-volatility scenarios.
While there remains potential for refining the Naïve models, the current results validate the advantages of the proposed adjustments. Notably, even a Naïve scores model demonstrates promise as a foundational tool for forecasting prospects, showcasing the adaptability and robustness of the approach.
The primary contribution of this paper lies not in detailing machine learning model implementations but in the introduction of a novel probability measure that innovatively connects sentiment analysis with market dynamics.
Beyond validating improvements in financial forecasting, these results pave the way for broader theoretical applications. In the subsequent section 4, we introduce the hype-adjusted probability measure designed to quantify and integrate market hype, thereby extending the framework beyond traditional sentiment analysis.
5.2 Future Work
Building upon the introduction of the hype-adjusted probability measure $\mathbb{P}^{a}$, our future work aims to extend this concept into a more comprehensive probability theory framework. By leveraging insights from the validation and calibration methods of the Mathematical Model lacking close-form solutions in recent work [cao2024theoreticalempiricalvalidationheston], we plan to develop an algorithm that computes the Radon-Nikodym derivative or Girsanov theorem set function and optimize constant and time-dependent parameters based on the forecast output of various models.
This approach will refine the sentiment model by systematically addressing data biases, such as political leanings or disproportionate media attention given to certain stocks (e.g., Nvidia receiving excessive coverage compared to its sector weight). The algorithm will also incorporate parameters like variance and precision in the loss function to improve prediction accuracy, assuming identical training models and data dimensions.
Future work will also explore integrating sentiment-guided adversarial learning frameworks, such as those presented in [sentiment-guided2023], to enhance the ability of the model to adapt to dynamic market conditions. By leveraging these advanced techniques, we aim to better capture the nuanced interplay between market hype and fundamental indicators, improving the generalizability of the hype-adjusted probability measure.
By aligning sentiment scores with security prices as time approaches maturity, this method will create a more objective measure to correct biases from the data sources mentioned earlier. The goal is to ensure that the sentiment-driven predictions are more reflective of market fundamentals, especially when sentiment diverges significantly from the stock’s actual influence in the sector.
Additionally, future research will explore the incorporation of further indicator functions into the model. For instance, introducing non-zero thresholds, $\mathbbm{1}_{|Sent_{d}-Sent_{d-1}|>c}$, could enhance the model by accounting for market behavior where significant impacts occur only when the sentiment change exceeds a certain threshold, $c$. This would provide a more nuanced understanding of how sharp sentiment shifts influence volatility and pricing, reinforcing the utility of the hype-adjusted probability measure $\mathbb{P}^{a}$ in forecasting financial markets.
Further analysis of bias or subjectivity could also be conducted using tools like the Eikon Data API, as discussed earlier in the paper. This will help deepen our understanding of the interplay between sentiment, volatility, and pricing, and further validate the efficacy of the hype-adjusted probability measure in real-world applications.
Our next paper is aiming to derive and formalize the adjusted volatility and option pricing proposals in section 4.2.3
6 Conclusion
In this paper, we presented an improved NLP approach for forecasting stock return and volatility based on the introduction of the new sentiment equation, the “3 weights", and the hype-adjusted probability measure, $\mathbb{P}^{a}$.
Our findings provide quantitative support for the market adage "buy on rumors, sell on news," as we confirm a stronger correlation between negative news sentiment and market volatility, as well as price declines. This highlights the asymmetrical influence of negative sentiment on the market compared to positive sentiment.
Our model evaluates text tokens with greater precision by integrating historical data, effectively capturing the influence of correlations and weights across subsets. A key advantage is its low sensitivity to the precision of the initial token scoring method, making it robust across varied data sources. Figure 8 demonstrates the enhanced performance of our sentiment model, showing higher accuracy in predicting market responses compared to baseline models.
Most importantly, we introduced the concept of a new probability measure, the hype-adjusted probability measure, $\mathbb{P}^{a}$, and outlined potential avenues for future research, including deriving new option pricing formulas, refining sentiment weighting parameters, and expanding applications within and beyond market forecasting. [^1] We are over the limit of our knowledge.
7 Acknowledgements
Zheng Cao would like to express his deepest gratitude to his mentor Professor Hélyette Geman for her invaluable teaching and supervision. Her guidance has been instrumental in shaping his understanding and approach. The term “hype-adjusted probability measure," $\mathbb{P}^{a}$, which she proposed, is both brilliant and inspiring.
We also wish to acknowledge the support and advice from Professor Zhen-Qing Chen and Jason Xie of the University of Washington, Wenyu Du of Yale University, and Professor James Schmidt, Xinyu Chang, and Kai Chen of Johns Hopkins University. Additionally, we appreciate Jieruo He of Johns Hopkins University’s assistance with data collection and constructing the sentiment score equation 19.
Lastly, we want to express our heartfelt gratitude for the insightful comments and constructive suggestions of the reviewers.
[…]
Appendix A Appendix A: Ticker News Weight Table
For the table below, Close Price refers to the price per share at market close in U.S. dollars; Capital is reported in millions of dollars; Capital Weight % represents the company’s market capitalization as a percentage of the entire sector; and News Weight % indicates the proportion of news coverage the company receives relative to the whole sector.
Ticker Name Close Price Capital Capital Weight % News Weight %
NVDA.OQ 131.38 1120867.0 8.66 24.52
INTC.OQ 34.59 524362.2 4.05 12.49
AMD.OQ 177.1 992494.8 7.67 6.11
TSM.N 184.52 528702.9 4.08 5.96
MU.OQ 131.14 490637.9 3.79 5.32
QCOM.OQ 207.12 752063.7 5.81 4.82
AVGO.OQ 1733.31 1220293.0 9.43 4.52
AMAT.OQ 251.47 866274.3 6.69 3.27
ASML.OQ 1059.97 477667.0 3.69 3.25
STM.N 41.51 137307.8 1.06 2.71
ON.OQ 73.48 394591.1 3.05 2.23
MRVL.OQ 73.84 501852.5 3.88 2.15
MCHP.OQ 92.34 444145.9 3.43 2.03
WOLF.N 23.05 35874.98 0.28 1.83
NXPI.OQ 274.91 472496.2 3.65 1.78
TXN.OQ 200.16 480016.7 3.71 1.77
ADI.OQ 232.01 462719.8 3.57 1.65
LRCX.OQ 1112.55 558002.5 4.31 1.54
UMC.N 8.59 78866.05 0.61 1.25
KLAC.OQ 874.9 538704.0 4.16 1.22
ASX.N 11.96 96245.99 0.74 1.21
QRVO.OQ 119.69 142502.9 1.10 1.15
TER.OQ 153.48 295777.2 2.28 0.97
MKSI.OQ 135.06 113224.3 0.87 0.95
SWKS.OQ 106.41 213374.7 1.65 0.94
ACLS.OQ 151.06 60574.45 0.47 0.91
ENTG.OQ 140.23 263961.8 2.04 0.88
LSCC.OQ 59.75 101910.3 0.79 0.87
MPWR.OQ 846.2 493544.3 3.81 0.76
RMBS.OQ 63.88 85799.32 0.66 0.72
Note, this table is collected from LSEG, on July 17, 2024[LSEG]. This ticker weight table is adjusted by removing 2 values: CME E-MINI S&P500-TECHNOLOGY SECTOR INDEX FUTURE SEP 2024, $0.15247\%$ and CME INDEX and OPTIONS MARKET E-MINI RUSSELL 2000 INDEX FUTURE SEP 2024, $0.15247\%$.